OpenStudy (anonymous):

Help with slope intercept from? Write the equation in slope-intercept from. Then find the slope an y-intercept of each line. 1. -1/2x - y = 3/4 2. 9x - 2y = 10 I don't really understand how to answer these problems, i'd appreciate the help. I under stand I need y on one side by itself, I just understand how.

OpenStudy (anonymous):

so slope intercept form is just y=mx+b

OpenStudy (anonymous):

so you just need to rearrange these functions so you get the form y=

OpenStudy (anonymous):

how do i rearrange it, that what i don't get

OpenStudy (anonymous):

so lets walk through the first one -1/2x - y = 3/4

OpenStudy (anonymous):

k

OpenStudy (anonymous):

would i add -1/2x to each side?

OpenStudy (anonymous):

close, you would add 1/2x to each side, to get rid of that term on the left,

OpenStudy (anonymous):

okay so it'd be y=1/2x+3/4 right? Then how would i find slope and y intercept?

OpenStudy (anonymous):

The slope is always "mx" and the y-intercept is always "b"

OpenStudy (anonymous):

ohhh i get it so yintercpt would be 3/4?

OpenStudy (anonymous):

what about slope?

OpenStudy (anonymous):

The slope is 1/2 because it is the number in front of the x

OpenStudy (anonymous):

close, you get -y=1/2x+3/4, but then you want y to be positive so you need to multiply the right side by-1, to get y=-1/2x-3/4

OpenStudy (anonymous):

so the slope is -1/2 and the y intercept is -3/4

OpenStudy (anonymous):

@ramiva is right, sorry I didn't see the negative sign in front of the y.

OpenStudy (anonymous):

Do you always want the y to be positive? And thank you (both) i think I get it now :)

OpenStudy (anonymous):

The y always has to be positive.

OpenStudy (anonymous):

yes you want the y to be positive in your final answer

OpenStudy (anonymous):

okay :)